how many calories were in the peanut if the following data are collected? mass of peanut burned= 0.25g volume of water heated = 100mL temperature change = 20'C specific heat of water = 1.00 cal/g'C
1
Expert's answer
2020-06-12T07:27:31-0400

The density of water is 0.998 g/mL. Therefore the mass = 100 x 0.998 = 99.8 g.

q = cm(T2 - T1) = 1.00cal/goC x 99.8g x 20oC = 1996 cal

Answer: 1996 cal or 1.996 Cal